# Combining Evidence Across Filtrations

### v4

@@ 6.1 Known characterization of adjusters for test supermartingales
We begin by reviewing the notion of “adjusters for test supermartingales.” This appears in each of the aforementioned early works, although the work of Dawid et al., 2011b is the most relevant here. Their work shows that an increasing function $\mathsf{A}:[1,\infty]\to[0,\infty]$ is an adjuster in the sense of definition (11) if and only if, for any probability distribution $P$ and for any test supermartingale $(M_{t})_{t\geq 0}$ for $P$ in some filtration $\mathbb{G}$, there exists another test supermartingale $(M_{t}^{\prime})_{t\geq 0}$ for $P$ in $\mathbb{G}$ that upper-bounds the adjusted running maximum process $(\mathsf{A}(M_{t}^{*}))_{t\geq 0}$:
@@
Dawid et al., 2011b ; Shafer et al., (2011) use this characterization as the definition and prove its equivalence to our definition (11). This alternative definition is now in terms of test supermartingales in some (arbitrary) probability and filtration, whereas definition (11) was purely analytic. It also does not describe whether the analogous characterization for e-processes would be equivalent, or how it relates to the fact that adjusters can perform $\mathfrak{e}$-lifting across filtrations.
@@ 6.2 The main characterization theorem
We now introduce the main result of this section, synthesizing the two characterizations of adjusters with Theorem 2. The takeaway is that all notions of adjusters coincide, including their characterization as $\mathfrak{e}$-lifters, as long as $\mathsf{A}$ is an increasing, deterministic function applied to the running maximum of an e-process.
@@
The proof of the latter statement leverages a technical lemma from Ramdas et al., (2020) that connects test supermartingales and e-processes. Given these two implications, we can establish that all five conditions are equivalent, via a $\Rightarrow$ e $\Rightarrow$ d $\Rightarrow$ 27 $\Rightarrow$ b $\Leftrightarrow$ a. The equivalence implies that admissibility according to any characterization must coincide with each other.
Theorem 3 establishes an extent to which the use of an adjuster is necessary for lifting evidence: the implication e $\Rightarrow$ a says that, if $f$ is some increasing function that maps the running maximum of any e-process in $\mathbb{G}$ into an e-process in any $\mathbb{F}\supseteq\mathbb{G}$, for an arbitrary $\mathcal{P}$, then $f$ must be an adjuster. Further, even if we found an increasing function $f$ such that, for any $\mathcal{P}$, its application to the running maximum of any e-process in $\mathbb{G}$ is an e-process in some $\mathbb{F}\supseteq\mathbb{G}$, then we can still be assured that $f$ is an adjuster because $f$ then satisfies condition d. Put together, we may conclude that, when it comes to mapping the running maximum of arbitrary e-processes to a valid e-process in some finer filtration, using an adjuster is necessary.
@@
We close with the note that the characterization of adjusters and $\mathfrak{e}$-lifters in Theorem 3 is specific to (increasing) functions that apply to the running maximum of e-processes. Thus, in theory, the theorem does not rule out the existence of an $\mathfrak{e}$-lifter that depends on, say, $\mathfrak{e}_{t}$ but not $\mathfrak{e}_{t}^{*}$, or other functions with entirely different forms. In Section E.2, we explore a class of generalized adjusters (Dawid et al., 2011a, ) that take both $\mathfrak{e}_{t}^{*}$ and $\mathfrak{e}_{t}$ as inputs, and we show how these are not $\mathfrak{e}$-lifters in general. There can also exist $\mathfrak{e}$-lifters for specific testing problems and for e-processes in particular filtrations, although their utility will be limited to such settings. Identifying alternate forms of $\mathfrak{e}$-lifters is left as future work.
@@ 7 Discussion
The $\mathfrak{e}$-lifting result establishes the answer to a theoretical question that, to the best of our knowledge, was not previously known in its generality. To motivate the question, we recall that when it comes to test martingales (not e-processes), there are testing problems for which there exist no powerful test martingales in the data filtration while there do exist powerful test martingales in a coarser filtration, such as Vovk, 2021b ’s conformal test martingale for randomness and Henzi and Law, (2024)’s test martingale for independence. We may ask the analogous question for e-processes:
@@
At a first glance, it may appear that the coarseness of the base filtration $\mathbb{G}$, relative to the data filtration $\mathbb{F}$, must somehow affect the power of the adjusted e-process. However, because the adjuster acts as a one-time, uniform fee for achieving anytime-validity in the finest “relevant” filtration $\mathbb{F}$, adjustment itself is independent of how coarse $\mathbb{G}$ is. On the other hand, the coarseness of $\mathbb{G}$ can affect the power of the base e-process $(\mathfrak{e}_{t})_{t\geq 0}$ in different ways. As we already saw, coarsening the filtration by an “appropriate” amount is a generally effective strategy for designing powerful test martingales for highly composite nulls (e.g., exchangeability and independence). Yet, if $\mathbb{G}$ is too coarse, e.g., if it contains no information about the data, then any e-process in $\mathbb{G}$ would have to be powerless.
@@
Since adjusters sacrifice some amount of evidence, we may consider leveraging external randomization to recover the loss in statistical power. In Section J, we explore this idea in two ways; here, we briefly summarize our findings. The first approach is to borrow ideas from Ramdas and Manole, (2024), who utilize an independent $\mathsf{Unif}[0,1]$ random variable $U$ and use the threshold of $U/\alpha$ for testing at a stopping time. This “lift-then-randomize” approach is straightforwardly applicable to the adjusted e-process. The second approach is to build upon the ideas of Ignatiadis et al., (2024), who develop a randomized e-to-p calibration method for e-values. However, we find that the resulting “randomize-then-lift” approach does not generally yield anytime-valid procedures when applied to $\mathfrak{e}$-lifting.
@@ Acknowledgements
AR acknowledges funding from NSF IIS-2229881 and NSF DMS-2310718. The authors thank the anonymous reviewers for their valuable feedback. YJC acknowledges Jan Hannig for suggesting the idea of using conditional expectations in Section 7.
@@ B.1 Problem setup
We remark that any nontrivial e-process for the IID null must be a nontrivial e-process for sequentially testing whether the data forms an exchangeable sequence (Ramdas et al.,, 2022). Recall that a sequence of random variables $(X_{t})_{t\geq 1}$ is exchangeable if, for any $t\geq 1$ and any permutation $\sigma$ over $\{1,\dotsc,t\}$, $(X_{1},\dotsc,X_{t})$ has the same distribution as $(X_{\sigma(1)},\dotsc,X_{\sigma(t)})$. Then, we may equivalently consider the following null:
@@ B.2 The universal inference e-process
Ramdas et al., (2022) derive an e-process for $\mathcal{H}_{0}^{\mathsf{exch}}$ in the data filtration $\mathbb{F}$ that is powerful against Markovian alternatives by leveraging universal inference (UI; Wasserman et al.,, 2020). The high-level idea of UI is to compute a generalized likelihood ratio, whose denominator is the maximum likelihood under the null and the numerator is a mixture (Robbins,, 1970) over a class of alternatives. For the denominator, it suffices to use the maximum likelihood under the IID family $\mathcal{H}_{0}^{\mathsf{IID}}=\{\mathsf{Ber}(p)^{\infty}:p\in[0,1]\}$, whose convex hull is precisely $\mathcal{H}_{0}^{\mathsf{exch}}$ due to de Finetti, (1931)’s theorem, and then invoke Ramdas et al., (2020, proposition 15)’s result that says any e-process for a family of distributions (e.g., $\mathcal{H}_{0}^{\mathsf{IID}}$) is also an e-process for its convex hull ($\mathcal{H}_{0}^{\mathsf{exch}}$). For the numerator, Ramdas et al., (2022) choose a tractable mixture distribution over all first-order Markovian alternatives, mixing over their two parameters $p_{0\to 1}$ and $p_{1\to 1}$, where $p_{j\to k}$ denotes the transition probability from state $j$ to state $k$ for $j,k\in\{0,1\}$.
@@ B.3 Conformal test martingales and the “simple jumper” algorithm
Vovk et al., (2003); Vovk, 2021b introduce a sequential testing procedure for exchangeability entirely based on conformal p-values. At each time $t$, define the nonconformity score $\alpha_{t}=\mathcal{A}(X_{t};X_{1},\dotsc,X_{t})\in\mathbb{R}$. For general data types, $\mathcal{A}$ can be any measure for the “strangeness” of $X_{t}$ alongside the observed data $X_{1},\dotsc,X_{t}$; for our binary case, we follow the prior works and use the identity function, that is, $\alpha_{t}=X_{t}$ for each $t\geq 1$ (Vovk, 2021a, ). Then, the conformal p-value $s_{t}$ at time $t$ is defined using the nonconformity scores $\alpha_{1},\dotsc,\alpha_{t}$ as follows:
@@
where $f_{k}(t)=t^{-1}\cdot\#\{i\leq t:X_{i}=k\}$ denotes the empirical frequency of seeing $k\in\{0,1\}$ up to time $t$, and $U_{t}\sim\mathsf{Unif}[0,1]$ is an IID Uniform random variable that is independent of $\mathbb{F}$. Each $s_{t}$ captures the “strangeness” of $X_{t}$ relative to the summary statistics (the empirical frequencies) of the data up to time $t$, along with external randomization. Equation (31) is a general definition for conformal p-values, while equation (32) represents a simplification assuming that the nonconformity score is the identity function and that the data is binary. By Vovk et al., (2003); Vovk et al., 2022a , the conformal p-values $(s_{t})_{t\geq 1}$ are IID with $\mathsf{Unif}[0,1]$ under the null hypothesis $\mathcal{H}_{0}^{\mathsf{IID}}$. Despite its name, the conformal p-value is not a p-value in the traditional sense, as it can approach either 0 or 1 under changepoint alternatives.
@@
In practice, we may “hedge our bets” across different values of $\lambda$, by taking the weighted average across (say) $\lambda\in\{-1,0,1\}$, and further “rebalance” a fraction of our portfolio at each round. These strategies yield Vovk et al., (2021)’s simple jumper in its full form:
@@
Empirically speaking, when we instead use $w_{-1}=w_{1}=1/2$ ($w_{0}=0$), with $\epsilon=0.01$, the mean of stopped e-values (as in right plot of Figure 1) also exceeds 1 when the stopping time is the first time when $k$ consecutive zeros or ones are observed, although the standard error over repeated runs is substantially larger. We find that Vovk, 2021b ’s original variant, with $w_{-1}=w_{0}=w_{1}=1/3$ and $\epsilon=0.01$, does not empirically violate the stopping time validity in the data filtration $\mathbb{F}$, despite not having $\mathbb{F}$-anytime-validity in theory. For the simulations under the alternative hypothesis (Figures 5 and 9), and for the real data experiments (Section 4.2), we use the original version with $w_{-1}=w_{0}=w_{1}=1/3$ and $\epsilon=0.01$.
In this paper, the specific choice of the simple jumper conformal test martingale is made solely for the clarity of exposition, and the naïve simple jumper algorithm is by no means the optimal version of the conformal test martingale. There exist other variants of conformal test martingales, including ones powerful against other types of alternatives; see Vovk et al., 2022a (, section 9). Yet, as referenced in the main text, even when combining different instantiations of conformal test martingales, the same issue of not having validity at data-dependent stopping times would still arise.
@@ Appendix C Additional example: sequentially testing a scale-invariant Gaussian mean
Here, we give another concrete example of an e-process defined in a coarser filtration $\mathbb{G}\subsetneq\mathbb{F}$ that is not valid at a $\mathbb{F}$-stopping time. This particular construction is due to Hendriksen et al., (2021); Pérez-Ortiz et al., (2024); we follow the exposition of the latter paper. The e-process described below can be combined with another e-process that is valid in $\mathbb{F}$, such as Wang and Ramdas, 2025a ’s e-process based on UI (Wasserman et al.,, 2020).
Example 4 (Sequentially testing a scale-invariant Gaussian mean; Hendriksen et al., (2021); Pérez-Ortiz et al., (2024)).
@@
Intuitively, $\mathbb{G}$ removes from $\mathbb{F}$ the scale information about the initial data point $X_{1}$ and retains only its sign. Then, the following ratio of likelihood mixtures (Cox,, 1952; Lai,, 1976),
@@
is an e-process for $\mathcal{H}_{0}$ in $\mathbb{G}$ (Pérez-Ortiz et al.,, 2024, proposition 1.3), where $\hat{\mu}_{t}$ and $\hat{\sigma}_{t}^{2}$ are the empirical mean and variance of $X_{1},\dotsc,X_{t}$. (Here, we sidestep group-theoretic notions in the original paper that are not central to our discussion.) Under the null, the expectation of $\mathfrak{e}^{\mathsf{MaxInv}}$ at any $\mathbb{G}$-stopping time $\tau^{\mathbb{G}}$ is upper-bounded by 1.
However, Pérez-Ortiz et al., (2024, app. B) also notes that $\mathfrak{e}^{\mathsf{MaxInv}}$ is not upper-bounded by 1 at $\mathbb{F}$-stopping times in expectation. Define $\tau^{\mathbb{F}}=1+\mathds{1}\left(\left\lvert X_{1}\right\rvert\in[a,b]\right)$ for some constants $0<a<b$, meaning that we stop immediately if the absolute value of $X_{1}$ is outside a range, and if not we stop after observing the second data point. While $\tau^{\mathbb{F}}$ is clearly an $\mathbb{F}$-stopping time, it is not adapted in $\mathbb{G}$ because it depends on the scale of $X_{1}$. For the values $\delta_{0}=0$, $a\approx 0.44$, $b\approx 1.70$, and a suitable choice of mixture over $\mathcal{H}_{1}$, numerical calculations show that
@@
For this problem, we can also consider combining different e-processes in an analogous manner to the case of testing randomness (Example 1). For instance, Wang and Ramdas, 2025a (, theorem 3.2) develop a UI-based e-process for sequentially testing whether a Gaussian mean is zero under unknown variance (the “Gaussian $t$-test”), which coincides with testing the scale-invariant null $\mathcal{H}_{0}$ when $\delta_{0}=0$. (We note that, in the paper, they further discuss test martingales for the general scale-invariant null.) Analogous to Ramdas et al., (2022)’s UI e-process for testing randomness (30), and Wang and Ramdas, 2025a ’s e-process is valid in the data filtration $\mathbb{F}$. It has the following closed-form expression:
@@
It is known that, generally speaking, the running maximum of a test martingale or an e-process for some null hypothesis $\mathcal{P}$ is not a test martingale or an e-process for $\mathcal{P}$, respectively; see, e.g., Shafer et al., (2011, answer 1) and Ramdas et al., (2020, remark 16). To show a concrete example of when this occurs, we describe a case where we can show something even stronger: the expected running maximum under the null can tend to infinity. Then, we discuss how this represents a broad class of test martingales (and, often, e-processes).
D.1 Illustrative example: Sequentially testing a bounded IID mean
We begin with the problem of sequentially testing the mean of IID bounded random variables. Suppose that $X_{1},X_{2},\dotsc\sim P$ are IID random variables taking values in $[0,1]$. Let $\mathbb{F}=(\mathcal{F}_{t})_{t\geq 0}$ where $\mathcal{F}_{t}=\sigma(X_{1},\dotsc,X_{t})$. We shall further assume that each $X_{t}$ has a finite but non-zero variance, $\sigma^{2}\in(0,\infty)$. Denoting $\mathcal{P}^{\mu}$ as all possible distributions on $[0,1]$ whose mean is $\mu$, we can test whether $P\in\mathcal{P}^{\mu}$ by constructing an e-process for the null hypothesis $\mathcal{P}^{\mu}$.
This is a composite and nonparametric testing problem that has been studied extensively in the SAVI literature. For example, Waudby-Smith and Ramdas, (2024) design powerful test martingales for $\mathcal{P}^{\mu}$ of the form

### v5

@@ 6.1 Known characterization of adjusters for test supermartingales
We begin by reviewing the notion of “adjusters for test supermartingales.” This appears in each of the aforementioned early works, although the work of Dawid et al. (2011b) is the most relevant here. Their work shows that an increasing function $\mathsf{A}:[1,\infty]\to[0,\infty]$ is an adjuster in the sense of definition (11) if and only if, for any probability distribution $P$ and for any test supermartingale $(M_{t})_{t\geq 0}$ for $P$ in some filtration $\mathbb{G}$, there exists another test supermartingale $(M_{t}^{\prime})_{t\geq 0}$ for $P$ in $\mathbb{G}$ that upper-bounds the adjusted running maximum process $(\mathsf{A}(M_{t}^{*}))_{t\geq 0}$:
@@
Dawid et al. (2011b); Shafer et al. (2011) use this characterization as the definition and prove its equivalence to our definition (11). This alternative definition is now in terms of test supermartingales in some (arbitrary) probability and filtration, whereas definition (11) was purely analytic. It also does not describe whether the analogous characterization for e-processes would be equivalent, or how it relates to the fact that adjusters can perform $\mathfrak{e}$-lifting across filtrations.
@@ 6.2 The main characterization theorem
We now introduce the main result of this section, synthesizing the two existing characterizations of adjusters with Theorem 2. The takeaway is that all notions of adjusters coincide, including their characterization as $\mathfrak{e}$-lifters, as long as $\mathsf{A}$ is an increasing, deterministic function applied to the running maximum of an e-process.
@@
The proof of the latter statement leverages a technical lemma from Ramdas et al. (2020) that connects test supermartingales and e-processes. Given these two implications, we can establish that all five conditions are equivalent, via a $\Rightarrow$ e $\Rightarrow$ d $\Rightarrow$ 27 $\Rightarrow$ b $\Leftrightarrow$ a. The equivalence implies that admissibility according to any characterization must coincide with each other.
Theorem 3 establishes an extent to which the use of an adjuster is necessary for lifting evidence. The implication e $\Rightarrow$ a says that, if $f$ is some increasing function that maps the running maximum of any e-process in $\mathbb{G}$ into an e-process in any finer filtration $\mathbb{F}\supseteq\mathbb{G}$, for an arbitrary $\mathcal{P}$, then $f$ must be an adjuster. Further, by d $\Rightarrow$ a, even if we only found an increasing function $f$ such that, for any $\mathcal{P}$, its application to the running maximum of any e-process in $\mathbb{G}$ is still an e-process in $\mathbb{G}$ (or in some $\mathbb{F}\supseteq\mathbb{G}$), then we can still be assured that $f$ is an adjuster. Put together, we may conclude: when it comes to mapping the running maximum of arbitrary e-processes to a valid e-process in some finer filtration, using an adjuster is necessary.
@@
We close with the note that the characterization of adjusters and $\mathfrak{e}$-lifters in Theorem 3 is specific to (increasing) functions that apply to the running maximum of e-processes. Thus, in theory, the theorem does not rule out the existence of an $\mathfrak{e}$-lifter that depends on, say, $\mathfrak{e}_{t}$ but not $\mathfrak{e}_{t}^{*}$, or other functions with entirely different forms. In Section E.2, we explore a class of generalized adjusters (Dawid et al., 2011a) that take both $\mathfrak{e}_{t}^{*}$ and $\mathfrak{e}_{t}$ as inputs, and we show how these are not $\mathfrak{e}$-lifters in general. There can also exist $\mathfrak{e}$-lifters for specific testing problems and for e-processes in particular filtrations, although their utility will be limited to such settings. Identifying alternate forms of $\mathfrak{e}$-lifters is left as future work.
@@ 7 Discussion
The $\mathfrak{e}$-lifting result establishes the answer to a theoretical question that, to the best of our knowledge, was not previously known in its generality. To motivate the question, we recall that when it comes to test martingales (not e-processes), there are testing problems for which there exist no powerful test martingales in the data filtration while there do exist powerful test martingales in a coarser filtration, such as Vovk (2021b)’s conformal test martingale for randomness and Henzi and Law (2024)’s test martingale for independence. We may ask the analogous question for e-processes:
@@
At a first glance, it may appear that the coarseness of the base filtration $\mathbb{G}$, relative to the data filtration $\mathbb{F}$, must somehow affect the power of the adjusted e-process. However, because the adjuster acts as a one-time, uniform fee for achieving anytime-validity in the finest “relevant” filtration $\mathbb{F}$, adjustment itself is independent of how coarse $\mathbb{G}$ is. On the other hand, the coarseness of $\mathbb{G}$ can affect the power of the base e-process $(\mathfrak{e}_{t})_{t\geq 0}$ in different ways. As we already saw, coarsening the filtration by an “appropriate” amount is a generally effective strategy for designing powerful test martingales for highly composite nulls (e.g., randomness and independence). Yet, if $\mathbb{G}$ is too coarse, e.g., if it contains no information about the data, then any e-process in $\mathbb{G}$ would have to be powerless.
@@
Since adjusters sacrifice some amount of evidence, we may consider leveraging external randomization to recover the loss in statistical power. In Section J, we explore this idea in two ways; here, we briefly summarize our findings. The first approach is to borrow ideas from Ramdas and Manole (2026), who utilize an independent $\mathsf{Unif}[0,1]$ random variable $U$ and use the threshold of $U/\alpha$ for testing at a stopping time. This “lift-then-randomize” approach is straightforwardly applicable to the adjusted e-process. The second approach is to build upon the ideas of Ignatiadis et al. (2024), who develop a randomized e-to-p calibration method for e-values. However, we find that the resulting “randomize-then-lift” approach does not generally yield anytime-valid procedures when applied to $\mathfrak{e}$-lifting.
@@ Acknowledgements
We thank the anonymous reviewers for their valuable feedback throughout the revision process. We acknowledge Jan Hannig for the idea of using conditional expectations in Section 7 and Martin Larsson for the proof sketch in Section D.2. A.R. acknowledges funding from NSF IIS-2229881 and NSF DMS-2310718.
@@ B.1 Problem setup
We remark that any nontrivial e-process for the IID null must be a nontrivial e-process for sequentially testing whether the data forms an exchangeable sequence (Ramdas et al., 2022). Recall that a sequence of random variables $(X_{t})_{t\geq 1}$ is exchangeable if, for any $t\geq 1$ and any permutation $\sigma$ over $\{1,\dotsc,t\}$, $(X_{1},\dotsc,X_{t})$ has the same distribution as $(X_{\sigma(1)},\dotsc,X_{\sigma(t)})$. Then, we may equivalently consider the following null:
@@ B.2 The universal inference e-process
Ramdas et al. (2022) derive an e-process for $\mathcal{H}_{0}^{\mathsf{exch}}$ in the data filtration $\mathbb{F}$ that is powerful against Markovian alternatives by leveraging universal inference (UI; Wasserman et al., 2020). The high-level idea of UI is to compute a generalized likelihood ratio, whose denominator is the maximum likelihood under the null and the numerator is a mixture (Robbins, 1970) over a class of alternatives. For the denominator, it suffices to use the maximum likelihood under the IID family $\mathcal{H}_{0}^{\mathsf{IID}}=\{\mathsf{Ber}(p)^{\infty}:p\in[0,1]\}$, whose convex hull is precisely $\mathcal{H}_{0}^{\mathsf{exch}}$ due to de Finetti (1931)’s theorem, and then invoke Ramdas et al. (2020, proposition 15)’s result that says any e-process for a family of distributions (e.g., $\mathcal{H}_{0}^{\mathsf{IID}}$) is also an e-process for its convex hull ($\mathcal{H}_{0}^{\mathsf{exch}}$). For the numerator, Ramdas et al. (2022) choose a tractable mixture distribution over all first-order Markovian alternatives, mixing over their two parameters $p_{0\to 1}$ and $p_{1\to 1}$, where $p_{j\to k}$ denotes the transition probability from state $j$ to state $k$ for $j,k\in\{0,1\}$.
@@ B.3 Conformal test martingales and the “simple jumper” algorithm
Vovk et al. (2003); Vovk (2021b) introduce a sequential testing procedure for exchangeability entirely based on conformal p-values. At each time $t$, define the nonconformity score $\alpha_{t}=\mathcal{A}(X_{t};X_{1},\dotsc,X_{t})\in\mathbb{R}$. For general data types, $\mathcal{A}$ can be any measure for the “strangeness” of $X_{t}$ alongside the observed data $X_{1},\dotsc,X_{t}$; for our binary case, we follow the prior works and use the identity function, that is, $\alpha_{t}=X_{t}$ for each $t\geq 1$ (Vovk, 2021a). Then, the conformal p-value $s_{t}$ at time $t$ is defined using the nonconformity scores $\alpha_{1},\dotsc,\alpha_{t}$ as follows:
@@
where $f_{k}(t)=t^{-1}\cdot\#\{i\leq t:X_{i}=k\}$ denotes the empirical frequency of seeing $k\in\{0,1\}$ up to time $t$, and $U_{t}\sim\mathsf{Unif}[0,1]$ is an IID Uniform random variable that is independent of $\mathbb{F}$. Each $s_{t}$ captures the “strangeness” of $X_{t}$ relative to the summary statistics (the empirical frequencies) of the data up to time $t$, along with external randomization. Equation (31) is a general definition for conformal p-values, while equation (32) represents a simplification assuming that the nonconformity score is the identity function and that the data is binary. By Vovk et al. (2003, 2005), the conformal p-values $(s_{t})_{t\geq 1}$ are IID with $\mathsf{Unif}[0,1]$ under the null hypothesis $\mathcal{H}_{0}^{\mathsf{IID}}$. Despite its name, the conformal p-value is not a p-value in the traditional sense, as it can approach either 0 or 1 under changepoint alternatives.
@@
In practice, we may “hedge our bets” across different values of $\lambda$ by taking the weighted average across (say) $\lambda\in\{-1,0,1\}$, and further “rebalance” a fraction of our portfolio at each round. These strategies yield Vovk et al. (2021)’s simple jumper in its full form:
@@
Empirically speaking, when we instead use $w_{-1}=w_{1}=1/2$ ($w_{0}=0$), with $\epsilon=0.01$, the mean of stopped e-values (as in right plot of Figure 1) also exceeds 1 when the stopping time is the first time when $k$ consecutive zeros or ones are observed, although the standard error over repeated runs is substantially larger. We find that Vovk (2021b)’s original variant, with $w_{-1}=w_{0}=w_{1}=1/3$ and $\epsilon=0.01$, does not empirically violate the stopping time validity in the data filtration $\mathbb{F}$, despite not having $\mathbb{F}$-anytime-validity in theory. For the simulations under the alternative hypothesis (Figures 5 and 9), and for the real data experiments (Section 4.2), we use the original version with $w_{-1}=w_{0}=w_{1}=1/3$ and $\epsilon=0.01$.
In this paper, the specific choice of the simple jumper conformal test martingale is made solely for the clarity of exposition, and the naïve simple jumper algorithm is by no means the optimal version of the conformal test martingale. There exist other variants of conformal test martingales, including ones powerful against other types of alternatives; see Vovk et al. (2005, section 9). Yet, as referenced in the main text, even when combining different instantiations of conformal test martingales, the same issue of not having validity at data-dependent stopping times would still arise.
@@ Appendix C Additional example: sequentially testing a scale-invariant Gaussian mean
Here, we give another concrete example of an e-process defined in a coarser filtration $\mathbb{G}\subsetneq\mathbb{F}$ that is not valid at a $\mathbb{F}$-stopping time. This particular construction is due to Hendriksen et al. (2021); Pérez-Ortiz et al. (2024); we follow the exposition of the latter paper. The e-process described below can be combined with another e-process that is valid in $\mathbb{F}$, such as Wang and Ramdas (2025a)’s e-process based on UI (Wasserman et al., 2020).
Example 4 (Sequentially testing a scale-invariant Gaussian mean; Hendriksen et al. (2021); Pérez-Ortiz et al. (2024)).
@@
Intuitively, $\mathbb{G}$ removes from $\mathbb{F}$ the scale information about the initial data point $X_{1}$ and retains only its sign. Then, the following ratio of likelihood mixtures (Cox, 1952; Lai, 1976),
@@
is an e-process for $\mathcal{H}_{0}$ in $\mathbb{G}$ (Pérez-Ortiz et al., 2024, proposition 1.3), where $\hat{\mu}_{t}$ and $\hat{\sigma}_{t}^{2}$ are the empirical mean and variance of $X_{1},\dotsc,X_{t}$. (Here, we sidestep group-theoretic notions in the original paper that are not central to our discussion.) Under the null, the expectation of $\mathfrak{e}^{\mathsf{MaxInv}}$ at any $\mathbb{G}$-stopping time $\tau^{\mathbb{G}}$ is upper-bounded by 1.
However, Pérez-Ortiz et al. (2024, app. B) also notes that $\mathfrak{e}^{\mathsf{MaxInv}}$ is not upper-bounded by 1 at $\mathbb{F}$-stopping times in expectation. Define $\tau^{\mathbb{F}}=1+\mathds{1}\left(\left\lvert X_{1}\right\rvert\in[a,b]\right)$ for some constants $0<a<b$, meaning that we stop immediately if the absolute value of $X_{1}$ is outside a range, and if not we stop after observing the second data point. While $\tau^{\mathbb{F}}$ is clearly an $\mathbb{F}$-stopping time, it is not adapted in $\mathbb{G}$ because it depends on the scale of $X_{1}$. For the values $\delta_{0}=0$, $a\approx 0.44$, $b\approx 1.70$, and a suitable choice of mixture over $\mathcal{H}_{1}$, numerical calculations show that
@@
For this problem, we can also consider combining different e-processes in an analogous manner to the case of testing randomness (Example 1). For instance, Wang and Ramdas (2025a, theorem 3.2) develop a UI-based e-process for sequentially testing whether a Gaussian mean is zero under unknown variance (the “Gaussian $t$-test”), which coincides with testing the scale-invariant null $\mathcal{H}_{0}$ when $\delta_{0}=0$. (We note that, in the paper, they further discuss test martingales for the general scale-invariant null.) Analogous to Ramdas et al. (2022)’s UI e-process for testing randomness (30), and Wang and Ramdas (2025a)’s e-process is valid in the data filtration $\mathbb{F}$. It has the following closed-form expression:
@@
It is known that, generally speaking, the running maximum of a test martingale or an e-process for some null hypothesis $\mathcal{P}$ is not a test martingale or an e-process for $\mathcal{P}$, respectively; see, e.g., Shafer et al. (2011, answer 1) and Ramdas et al. (2020, remark 16). To show a concrete example of when this occurs, we describe a case where we can show something even stronger: the expected running maximum under the null can tend to infinity. Then, we discuss how this represents a broad class of test martingales (and, often, e-processes).
D.1 Illustrative case: Sequentially testing a bounded IID mean
For clarity, we first focus on the popular problem of sequentially testing the mean of IID bounded random variables. Suppose that $X_{t}\sim P,\,\forall t$, are IID random variables taking values in $[0,1]$. Let $\mathbb{F}=(\mathcal{F}_{t})_{t\geq 0}$, $\mathcal{F}_{t}=\sigma(X_{1},\dotsc,X_{t})$. We assume that each $X_{t}$ has a finite but non-zero variance, $\sigma^{2}\in(0,\infty)$. Denoting $\mathcal{P}^{\mu}$ as all distributions on $[0,1]$ with mean $\mu\in(0,1)$, we want to test the null hypothesis $\mathcal{H}_{0}:P\in\mathcal{P}^{\mu}$ by constructing an e-process for $\mathcal{P}^{\mu}$.
This is a composite and nonparametric testing problem that has been studied extensively in the SAVI literature. For example, Waudby-Smith and Ramdas (2024) design powerful test martingales for $\mathcal{P}^{\mu}$ of the form